Torinto AXE Ltd is a United States company doing a financial plan for the following year. Currency Total Inflow Total Outflow Mexican Pesos (MXP) 15 000 000 9 000 000 Singapore dollars (S$) 7 000 000 9 000 000 New Zealand dollars (NZ$) 6 500 000 3 200 000 Canadian dollars (C$ 21 000 000 4 000 000 The spot rates and one-year forward rates as of today are: Currency Spot rate One-year forward rate MXP US$0.05 US$0.04 S$ US$0.73 US$0.71 NZ$ US$0.66 US$0.63 C$ US$0.76 US$0.79 Using the information given above, calculate the net exposure of each foreign currency in United States dollars


1
Expert's answer
2021-01-28T11:39:58-0500

Mexican Pesos (MXP): 15,000,000- 9,000,000=6,000,000

                       1 MXP = US$0.050

                       6,000,000 = US$ 300,000

                       300,000/0.04=7,500,000*0.05= 375,000

                       Net exposure= US$ 375,000

 

Singapore dollars(S$): 7,000,000-9,000,000= -2,000,000

                                   1 S$ = US$ 0.75

                                   S$ -2,000,000 = US$ -1,500,000

                                   -1,500,000/0.71= -2,112,676.06*0.73=1,542,253.53               

  Net exposure = 1,542,253.53  



New Zealand dollars (NZ$): 6,500,000- 3,200,000= 3,300,000

                                   NZ$ 1= US$ 0.72

                                   NZ$ 3,300,000 = 2,376,000

                                  2,376,000/0.63= 3,771,428.57* 0.66=2,489,142.86

                                  Net exposure= US$ = 2,489,142.86

           

 

Canadian dollars (C$) = 21,000,000- 4,000,000= 17,000,000

                                   C$ 1 = US$ 0.78

                                   C$ 17,000,000 = 13,260,000

13,260,000/0.79= 16,784,810.01*0.76= 12,756,455.70

                                   Net exposure = US$ 12,756,455.70